# SDPO: Segment-Level Direct Preference Optimization for Social Agents

### v1

@@ Abstract
Social agents powered by large language models (LLMs) can simulate human social behaviors but fall short in handling complex goal-oriented social dialogues. Direct Preference Optimization (DPO) has proven effective in aligning LLM behavior with human preferences across a variety of agent tasks. Existing DPO-based approaches for multi-turn interactions are divided into turn-level and session-level methods. The turn-level method is overly fine-grained, focusing exclusively on individual turns, while session-level methods are too coarse-grained, often introducing training noise. To address these limitations, we propose Segment-Level Direct Preference Optimization (SDPO), which focuses on specific key segments within interactions to optimize multi-turn agent behavior while minimizing training noise. Evaluations on the SOTOPIA benchmark demonstrate that SDPO-tuned agents consistently outperform both existing DPO-based methods and proprietary LLMs like GPT-4o, underscoring SDPO’s potential to advance the social intelligence of LLM-based agents. We release our code and data at this url.
@@ 1 Introduction
Figure 1: An overview of three alignment algorithms with varying granularities, illustrated using a SOTOPIA task as an example. Refer to caption represents the agent to be tested. Positive sessions achieve higher scores on the goal and relationship dimensions. For ease of presentation, we simplify the interaction content while preserving the original meaning.
@@
In response to these challenges, several methods have been developed to better align LLM behavior with human preferences in multi-turn interactions. These approaches offer promising strategies for improving social decision-making in LLMs. Specifically, we focus on Direct Preference Optimization (DPO)-based methods, which can be divided into two primary categories: turn-level DPO Rafailov et al. (2023), and session-level DPO approaches, such as ETO Song et al. (2024) and DMPO Shi et al. (2024). Turn-level DPO, for example, identifies a single conversational turn and uses a "positive-negative" pair of responses from that turn to optimize the model via a preference loss function. While turn-level DPO has demonstrated some effectiveness, its focus on individual turns limits its ability to model goal completion in goal-oriented social dialogues, where success often relies on high-quality interactions spanning multiple conversational turns.
To more effectively align agent behavior in multi-turn interactions, session-level alignment methods such as ETO and DMPO have been proposed. These methods extend the sampling scope from individual turns to entire sessions, constructing pairs of good and bad sessions and applying an adapted DPO loss for training. However, these methods still exhibit certain limitations due to their relatively coarse alignment granularity:
(i) Turns without errors in the negative sessions are still treated as bad outputs, introducing significant noise that can negatively affect the training process.
@@
To overcome the limitations of turn-level and session-level methods, we propose segment-level preference optimization, called SDPO, which aligns agent behavior in multi-turn interactions by optimizing turns within specific segments. Specifically, SDPO first identifies the erroneous turn in the negative session and then uses the interaction history preceding that turn to perform multiple samplings, thereby generating the positive session. Next, SDPO takes the first differing turn as the starting point, identifies the key segment from the positive session that contributes to a higher score, and forms data pairs by taking the corresponding segment with the same length from the negative session. Finally, an adapted DPO loss is calculated for the turns within the segments. We present an overview of three alignment algorithms at different granularities for social dialogues in Figure 1. Compared to turn-level DPO, SDPO aligns multiple interaction turns, making it more suitable for goal-oriented social dialogues. Our method also addresses the drawbacks of session-level algorithms:
(i) By calculating the loss only for the erroneous segments and their corresponding positive segments, training noise caused by non-erroneous turns is eliminated.
@@
We evaluate our approach on SOTOPIA Zhou et al. (2024), an open-ended and interactive goal-oriented benchmark for social intelligence, using both self-chat and interactions with other agents including GPT-4o and GPT-4o-mini, as interlocutors. Our results demonstrate that the SDPO-tuned agent consistently outperforms DPO, ETO, DMPO, and other robust baselines like GPT-4o, confirming the efficacy of segment-level alignment.
Essentially, turn-level and session-level alignments are special cases of segment-level alignment. When the segment length is 1, SDPO simplifies to DPO. When the segment encompasses the entire interaction from start to finish, SDPO transforms into ETO. Compared to existing methods, our approach is more flexible and general, allowing for the selection of the optimal alignment granularity based on the specific data. Moreover, in this paper, we primarily apply SDPO to enhance agents’ social intelligence. However, we believe our approach can be applied to other scenarios, further enhancing the capabilities of agents across various domains.
@@
  • We propose SDPO, a novel approach to align an agent’s behavior at the segment level in multi-turn interactions, supported by a rigorous theoretical framework.
  • We thoroughly evaluate our method on SOTOPIA, a simulated and interactive social benchmark, substantiating the efficacy of segment-level alignment.
  • We conduct an in-depth analysis of SDPO, validating its robustness across the variation in model output length, the efficiency of data utilization, the sources of data pairs, and the selection of segments.
@@ 2.1 SOTOPIA Environment
Unlike previous social benchmarks that primarily test through static QA formats Sap et al. (2019); Chen et al. (2024), SOTOPIA offers an interactive, open-ended, and realistic simulation environment, enabling a more precise assessment of agents’ social intelligence. A social task in SOTOPIA involves a scenario, two role profiles, and their private social goals to be achieved through interaction. The diverse combinations of scenarios and social goals encompass a broad spectrum of social interactions, such as negotiation, collaboration, and competition. SOTOPIA defines seven dimensions for evaluating social agents. We focus primarily on the "goal" (0-10, int) and "relationship" (-5 to 5, int), as GPT-4o’s ratings in these metrics closely align with human evaluations. SOTOPIA-$\pi$ Wang et al. (2024b) is a follow-up work that leverages GPT-4 to automatically construct a set of scenarios (completely non-overlapping with SOTOPIA), which serves as the training dataset for our study. Additionally, we restructure the prompt organization format of SOTOPIA to support multi-turn alignment, and the details are provided in Appendix A.
@@ 2.2 Task Formulation
Here, $y_{i}\sim\pi_{\theta}(\cdot|h_{i})$ represents the output generated by the LLM-based agent in round $i$ according to its policy $\pi_{\theta}$ with parameter $\theta$. On the other hand, $y_{i}^{{}^{\prime}}$ represents the output of the interlocutor, which is drawn from an unknown distribution. Based on this formulation, we present the DPO and ETO loss functions (ETO directly extends turn-level DPO to the session level without rigorous proof) in Appendix B.1 and B.2.
2.3 DMPO
DMPO introduces the state-action occupancy measure (SAOM) to help derive the loss of DPO in multi-turn interaction scenarios. The discounted SAOM $d^{\pi}(s,a)$ of a policy $\pi$ is as follows:
$d^{\pi}(s=s_{t},a=a_{t})=\gamma^{t}\cdot P(s_{0})\cdot\\ \prod_{k=0}^{t-1}\pi(a_{k}|s_{k})P(s_{k+1}|s_{k},a_{k}).$ (2)
Similar to DPO, DMPO defines the following RL objective based on $d^{\pi}$:
$\max_{\pi_{\theta}}\mathbb{E}_{(s,a)\sim d^{\pi_{\theta}}(s,a)}[r(s,a)]\\ -\beta\mathbb{D}_{KL}[d^{\pi_{\theta}}(s,a)||d^{\pi_{ref}}(s,a)],$ (3)
where $\pi_{ref}$ represents the reference policy. The reward function in DMPO takes the form:
$r(s,a)=\beta\log\frac{d^{\pi^{*}}(s,a)}{d^{\pi_{ref}}(s,a)}+\beta\log Z,$ (4)
where $Z$ is the partition function and remains constant for all $(s,a)$ pairs. Then apply the BT model as follows:
$p(\tau^{w}\succ\tau^{l}|s_{0})=\\ \sigma\left(\sum_{t=0}^{T_{w}-1}\gamma^{t}r(s_{t}^{w},a_{t}^{w})-\sum_{t=0}^{T_{l}-1}\gamma^{t}r(s_{t}^{l},a_{t}^{l})\right),$ (5)
where $\tau^{w}$ and $\tau^{l}$ represent the "win" and "lose" trajectories respectively, $T_{w},T_{l}$ denote the number of rounds in each. However, since $T^{w}\neq T^{l}$, the partition function $Z$ can not be canceled directly in Eq (5). To overcome this obstacle, DMPO introduces the length normalization and gets the final loss function (in the context of social dialogue):
A closer examination of Eq (5) reveals that the summation over the $(s,a)$ pairs should exclude $\gamma^{t}$, as it is already incorporated into $d^{\pi}(s,a)$. Furthermore, the length normalization may introduce additional assumptions that do not always hold. A detailed discussion of these issues are provided in Appendix B.3. Building on Eq. (5) without $\gamma^{t}$, we derive the SDPO loss in Section 3.3.
@@ 3.2 Preference Data Construction
Building high-quality segment-level preference data pairs is the core of our approach. On SOTPIA-$\pi$, our social agent engages in self-chat and interactions with GPT-4o. We set a threshold of 7 for the goal dimension, and all dialogues with a goal completion level below this threshold are considered potential negative samples. Given a negative session, the pipeline for generating positive data involves three steps, as illustrated in Figure 6.
Error Locaton Unlike scenarios with clear error definitions such as math, errors in social dialogues are a relatively ambiguous concept. In a negative session, if our agent’s utterance in a specific turn meets the following criteria: (1) the turn is critical for achieving the role’s goal, (2) there is still room for improvement in the goal completion or their relationship, we identify that turn as erroneous. The error location is performed by GPT-4o.
Positive Session Sampling After the error location, we sample 5 complete sessions based on the interaction history prior to that turn. Among these sessions, we select the one with the highest goal and relationship scores (with goal completion prioritized over relationship). If the goal or relationship score of the optimal session is higher than that of the negative sample, this session and the negative sample form a data pair; otherwise, the negative sample is discarded.
Segment Selection Once we obtain session-level data pairs, we provide both the positive and negative samples to GPT-4o, prompting it to select a segment from the positive sample. This segment should correspond to the part that contributes to the positive sample achieving higher goal and relationship scores. Subsequently, we extract a segment of the same length from the negative sample and pair it with the positive sample to form a segment-level data pair. This process aims to exclude turns, such as pleasantries, that are not directly related to achieving the goal.
The discussion of GPT-4o’s performance on error location and segment selection, along with the related prompts for these two steps, are provided in Appendix C.2 and C.3, respectively..
@@ 3.3 SDPO Loss
Session-level ETO and DMPO can not control the length of positive and negative sessions. To address this, DMPO utilizes the length normalization to eliminate $Z$ in Eq (5). Different from them, SDPO selects a segment from both the positive and negative sessions for optimization, allowing free control over their lengths. Specifically, after selecting a segment from the positive session using GPT-4o, SDPO selects a segment of equal length from the negative session. By ensuring the two segments are of equal length, we can directly eliminate $Z$, resulting in the following concise SDPO loss:
@@ 4 Experiments
Table 1: The performance of various methods on SOTOPIA across the goal and relationship dimensions. Additionally, SOTOPIA designates the more challenging portion of the dataset as the Hard subset, where SDPO also achieves the best results. Detailed results and discussion are presented in Appendix E.
@@
Table 2: The performance of different methods on SOTOPIA using Mistral.
@@ 4.1 Datasets
SOTPIA-$\pi$, used for training, includes a total of 410 scenarios: 100 scenarios for BC, with 10 role pairs per scenario, and 310 scenarios for alignment, with 8 role pairs per scenario. SOTOPIA, used for testing, includes 90 scenarios, each with 5 role pairs, resulting in a total of 450 tasks for self-chat and 900 tasks for non-self-chat.
@@ 4.2 Experimental Setup
Training We primarily use Llama-3.1-8B-Chat Dubey et al. (2024) as the base LLM to build the social agent. The maximum token limit is set to 4096, and AdamW optimization is employed for all training processes. During the SFT phase, the batch size is 32, the dropout rate is 0.2, and the learning rate is $1e^{-5}$ with $5\%$ warm-up ratio and a cosine decay schedule. For the training phase of SDPO, the batch size remains 32, $\beta$ in SDPO loss is 0.1, and the learning rate is $1e^{-6}$ with no warm-up but a cosine decay schedule. The statistics of SDPO training data are detailed in Appendix C.1.
SOTOPIA During the sampling of positive data, the temperature of the target agent is set to 1.0, while the other agent’s temperature is set to 0.7. For testing, we set the temperature of both interacting agents to 0.7. Though the temperature introduces randomness to the agents’ outputs, we find that the evaluation results remain numerically stable. Thus, we report the results based on a single test.

### v2

@@ Abstract
Social agents powered by large language models (LLMs) can simulate human social behaviors but fall short in handling complex social dialogues. Direct Preference Optimization (DPO) has proven effective in aligning LLM behavior with human preferences across various agent tasks. However, standard DPO focuses solely on individual turns, which limits its effectiveness in multi-turn social interactions. Several DPO-based multi-turn alignment methods with session-level data have shown potential in addressing this problem. While these methods consider multiple turns across entire sessions, they are often overly coarse-grained, introducing training noise, and lack robust theoretical support. To resolve these limitations, we propose Segment-Level Direct Preference Optimization (SDPO), which dynamically select key segments within interactions to optimize multi-turn agent behavior. SDPO minimizes training noise and is grounded in a rigorous theoretical framework. Evaluations on the SOTOPIA benchmark demonstrate that SDPO-tuned agents consistently outperform both existing DPO-based methods and proprietary LLMs like GPT-4o, underscoring SDPO’s potential to advance the social intelligence of LLM-based agents. We release our code and data at this url.
@@ 1 Introduction
In response to these challenges, several methods have been developed to better align LLM behavior with human preferences. These approaches offer promising strategies for improving social decision-making in LLMs. Specifically, we focus on Direct Preference Optimization (DPO)-based methods. Standard DPO Rafailov et al. (2023), identifies a single conversational turn and uses a ‘positive-negative’ pair of responses from that turn to optimize the model via a preference loss function. While DPO has demonstrated some effectiveness, its focus on individual turns limits its ability to model goal completion in goal-oriented social dialogues, where success often relies on high-quality interactions spanning multiple conversational turns.
To more effectively align agent behavior in multi-turn interactions, several multi-turn alignment methods including ETO Song et al. (2024) and DMPO Shi et al. (2024) have been proposed. These methods extend the sampling scope from individual turns to entire sessions, constructing pairs of good and bad sessions and applying an adapted DPO loss for training. We categorize these methods as session-level DPO. However, they exhibit limitations in both data granularity and theoretical foundations. From the data perspective, session-level DPOs suffer from the following drawbacks due to their relatively coarse alignment granularity:
(i) Turns without errors in negative sessions are also treated as bad outputs, introducing substantial noise that negatively affects the training process.
@@
From a theoretical perspective, in multi-turn scenarios, simply applying DPO fails to eliminate the partition function $Z$ Shi et al. (2024). ETO extends the DPO loss to multi-turn interactions but lacks a formal theoretical guarantee. DMPO incorporates SAOM theory and successfully converts $Z$ into a constant; however, due to the differing number of turns between positive and negative sessions, DMPO heuristically normalizes the length to eliminate $Z$ without a rigorous proof. A detailed theoretical analysis is provided in Appendix B.2.
To overcome the limitations of session-level DPO, we propose segment-level direct preference optimization (SDPO). Our approach shifts the sampling starting point backward and truncates the useless content at the end of sessions, thereby obtaining key segment pairs to refine granularity. Meanwhile, we ensure that the turn number of positive and negative segments is balanced, eliminating $Z$ and rigorously deriving the SDPO loss. Specifically, SDPO first identifies the erroneous turn in the negative session and then uses the interaction history preceding that turn to perform multiple samplings, thereby generating the positive session. Next, SDPO takes the first differing turn as the starting point, selects the key segment from the positive session that contributes to a higher score, and forms data pairs by taking the corresponding segment with the same length from the negative session. Finally, the SDPO loss is calculated for the turns within the segments. We provide an overview of three types of alignment algorithms for social dialogues in Figure 1. At the data level, our method can address the drawbacks of session-level DPO:
(i) By calculating the loss only for turns in negative and positive segments, training noise caused by non-erroneous turns is largely eliminated.
@@
In theory, due to the flexibility in segment selection, SDPO can control the number of turns in the positive and negative segments to ensure consistency, thereby eliminating $Z$ and yielding a simple yet rigorous SDPO loss.
We empirically evaluate our approach on SOTOPIA Zhou et al. (2024), an open-ended and interactive benchmark for social intelligence, using both self-chat and interactions with other agents including GPT-4o and GPT-4o-mini, as interlocutors. Our results demonstrate that the SDPO-tuned agent consistently outperforms DPO, ETO, DMPO, and other robust baselines like GPT-4o, confirming the efficacy of segment-level alignment.
Segment level is a more flexible and unified data granularity that dynamically selects the optimization scope for different data pairs, while also elegantly addressing the theoretical challenges of multi-turn alignment. In this paper, we primarily apply SDPO to enhance agents’ social intelligence. However, we believe our approach can be applied to other scenarios, further enhancing the capabilities of agents across various domains.
Figure 1: An overview of three alignment algorithms, illustrated using a SOTOPIA task as an example. Refer to caption represents the agent to be tested. A more detaild illustration is provided in Figure 6.
@@
  • We propose SDPO, a novel multi-turn alignment algorithm that provides a construction pipeline for segment-level preference data pairs, enabling more effective alignment granularity.
  • We identify theoretical limitations in existing multi-turn alignment methods and resolve them by leveraging the flexibility of the segment-level granularity, rigorously deriving a concise SDPO loss formula.
  • We evaluate our approach on SOTOPIA, a simulated and interactive social benchmark. The performance of SDPO, along with in-depth analysis demonstrates the efficacy and robustness of the segment-level alignment.
@@ 2.1 SOTOPIA Environment
Unlike previous social benchmarks that primarily test through static QA formats Sap et al. (2019); Chen et al. (2024), SOTOPIA offers an interactive, open-ended, and realistic simulation environment, enabling a more precise assessment of agents’ social intelligence. A social task in SOTOPIA involves a scenario, two role profiles, and their private social goals to be achieved through interaction. The diverse combinations of scenarios and social goals encompass a broad spectrum of social interactions, such as negotiation, collaboration, and competition. SOTOPIA defines seven dimensions for evaluating social agents. We focus primarily on the ‘goal’ (0-10, int) and ‘relationship’ (-5 to 5, int), as GPT-4o’s ratings in these metrics closely align with human evaluations. SOTOPIA-$\pi$ Wang et al. (2024b) is a follow-up work that leverages GPT-4 to automatically construct a set of scenarios (completely non-overlapping with SOTOPIA), which serves as the training dataset for our study. Additionally, we restructure the prompt organization format of SOTOPIA to support multi-turn alignment, and the details are provided in Appendix A.
@@ 2.2 Task Formulation
Here, $y_{i}\sim\pi_{\theta}(\cdot|h_{i})$ represents the output generated by the LLM-based agent in round $i$ according to its policy $\pi_{\theta}$ with parameter $\theta$. On the other hand, $y_{i}^{{}^{\prime}}$ represents the output of the interlocutor, which is drawn from an unknown distribution. Based on this formulation, we present the ETO and DMPO loss functions in Appendix B.1 and B.2.
2.3 Direct Preference Optimization
Rafailov et al. (2023) propose Direct Preference Optimization (DPO), a method leverages pairwise preference data to train policy models without relying on reinforcement learning Ouyang et al. (2022). In the context of social dialogue, we denote the number of the erroneous round as $e$, the DPO loss function is as follows:
where $y_{e}^{w},y_{e}^{l}\sim\pi_{\theta}(\cdot|h_{e})$ represent positive and negative output in the erroneous round respectively. However, due to its single-turn optimization nature, DPO is less suited for social dialogues. Extending DPO to multi-turn scenarios rigorously presents a challenge, which we will address in Section 3.3.
@@ 3.2 Preference Data Construction
Building high-quality segment-level preference data pairs is the core of our approach. On SOTPIA-$\pi$, our social agent engages in self-chat and interactions with GPT-4o. We set a threshold of 7 for the goal dimension, and all dialogues with a goal completion level below this threshold are considered potential negative samples. Given a negative session, the pipeline for generating segment-level data pairs involves three steps, as illustrated in Figure 2.
Error Location Unlike scenarios with clear error definitions such as math, errors in social dialogues are a relatively ambiguous concept. In a negative session, if our agent’s utterance in a specific turn meets the following criteria: (1) the turn is critical for achieving the role’s goal, (2) there is still room for improvement in the goal completion or their relationship, we identify that turn as erroneous. The error location is performed by GPT-4o, with the prompt provided in Appendix C.3
Positive Session Sampling After the error location, we sample 5 complete sessions based on the interaction history prior to that turn. Among these sessions, we select the one with the highest goal and relationship scores (priority given to goal). If the goal or relationship score of the optimal session is higher than that of the negative sample, this session and the negative session form a data pair; otherwise, the negative sample is discarded.
Segment Selection Once we obtain session-level data pairs, we provide both the positive and negative samples to GPT-4o, prompting it to select a segment from the positive sample. This segment should correspond to the part that contributes to the positive sample achieving higher goal and relationship scores. The prompt is also provided in Appendix C.3. Subsequently, we extract a segment of the same length from the negative sample and pair it with the positive sample to form a segment-level data pair. This process aims to exclude turns, such as pleasantries, that are not directly related to achieving the goal.
We evaluate GPT-4o’s performance in error location and segment selection, concluding that it effectively handles both tasks. Further details are provided in Appendix C.2.
@@ 3.3 SDPO Loss
We transfer the framework of DMPO to the dialogue domain and first introduce the state-action occupancy measure (SAOM). In this context, the interaction history $h$ serves as the state, while the agent’s output $y$ represents the action. The discounted SAOM $d^{\pi}(h,y)$ of a policy $\pi$ is as follows:
$d^{\pi}(h=h_{t},y=y_{t})=\gamma^{t}\cdot P(h_{0})\cdot\\ \prod_{k=0}^{t-1}\pi(y_{k}|h_{k})P(h_{k+1}|h_{k},y_{k}).$ (3)
The RL objective based on $d^{\pi}$ is as follows:
$\max_{\pi_{\theta}}\mathbb{E}_{(h,y)\sim d^{\pi_{\theta}}(h,y)}[r(h,y)]\\ -\beta\mathbb{D}_{KL}[d^{\pi_{\theta}}(h,y)||d^{\pi_{ref}}(h,y)],$ (4)
where $\pi_{ref}$ represents the reference policy, $\mathbb{D}_{KL}$ denotes the KL divergence. Following DPO, the optimal solution to the RL objective in Eq (4) is:
$d^{\pi^{*}}(h,y)=\frac{1}{Z}d^{\pi_{ref}}(h,y)\exp(\frac{1}{\beta}r(h,y)),$ (5)
where $\pi^{*}$ denotes the optimal policy, $Z$ is the partition function that normalizes the probability. As $d^{\pi}(h,y)$ is a function of $(h,y)$ pairs, normalizing it results in the partition functions $Z$ being independent of the current history $h$. Consequently, $Z$ remains constant for all $(h,y)$ pairs, marking a crucial step in their elimination. The reward function takes the form:
$r(h,y)=\beta\log\frac{d^{\pi^{*}}(h,y)}{d^{\pi_{ref}}(h,y)}+\beta\log Z.$ (6)
The Bradley-Terry (BT) model Bradley and Terry is then used to model the preference distribution. In this step, DMPO incorrectly duplicates the calculation of $\gamma$ in Eq. (3) and heuristically normalizes the length to eliminate $Z$ without a rigorous proof in subsequent steps. A detailed discussion of these issues are provided in Appendix B.2. Given preference data pairs, the correct application of the BT model is as follows:
$p(\tau^{w}\succ\tau^{l}|h_{0})=\\ \sigma\left(\sum_{t=0}^{T_{w}-1}r(h_{t}^{w},y_{t}^{w})-\sum_{t=0}^{T_{l}-1}r(h_{t}^{l},y_{t}^{l})\right),$ (7)
where $\tau^{w}$ and $\tau^{l}$ represent the ‘win’ and ‘lose’ samples respectively, $T_{w},T_{l}$ denote the number of rounds in each. However, session-level DPO can not control the length of positive and negative sessions, and since $T^{w}\neq T^{l}$ typically, the partition function $Z$ can not be canceled directly in Eq (7).
Different from them, SDPO selects a segment from both the positive and negative sessions for optimization, allowing free control over their lengths. By ensuring the two segments are of equal length, we can directly eliminate $Z$ in Eq (7). At the same time, by combining with Eq (3) to replace $d^{\pi}$, we obtain the following concise SDPO loss:
@@ 4 Experiments
Table 1: The performance of various methods on SOTOPIA across the goal and relationship dimensions. Additionally, SOTOPIA designates the more challenging portion of the dataset as the Hard subset, where SDPO also achieves the best results. Detailed results and discussion are presented in Appendix E.1.
@@
Table 2: The performance of different methods on SOTOPIA using Mistral-v0.3.
@@ 4.1 Datasets
SOTOPIA-$\pi$, used for training, includes a total of 410 scenarios: 100 scenarios for BC, with 10 role pairs per scenario, and 310 scenarios for alignment, with 8 role pairs per scenario. SOTOPIA, used for testing, includes 90 scenarios, each with 5 role pairs, resulting in a total of 450 tasks for self-chat and 900 tasks for non-self-chat.
@@ 4.2 Experimental Setup
Training We primarily use Llama-3.1-8B-Chat Dubey et al. (2024) as the base LLM to build the social agent. The maximum token limit is set to 4096, and AdamW optimizer is employed for all training processes. During the SFT phase, the batch size is 32, the dropout rate is 0.2, and the learning rate is $1e^{-5}$ with $5\%$ warm-up ratio and a cosine decay schedule. For the training phase of SDPO, the batch size remains 32, $\beta$ in SDPO loss is 0.1, and the learning rate is $1e^{-6}$ with no warm-up but a cosine decay schedule. The statistics of SDPO training data are detailed in Appendix C.1.
SOTOPIA During the sampling of positive data, the temperature of the target agent is set to 1.0, while the other agent’s temperature is set to 0.7. For testing, we set the temperature of both interacting agents to 0.7. Though the temperature introduces randomness to the agents’ outputs, we find that the evaluation results remain numerically stable. Thus, we report the results based on one single test.